# BPO: Towards Balanced Preference Optimization between Knowledge Breadth and Depth in Alignment

### v1

@@ Abstract
Reinforcement Learning with Human Feedback (RLHF) is the key to the success of large language models (LLMs) in recent years. In this work, we first introduce the concepts of knowledge breadth and knowledge depth, which measure the comprehensiveness and depth of an LLM or knowledge source respectively. We reveal that the imbalance in the number of prompts and responses can lead to a potential disparity in breadth and depth learning within alignment tuning datasets by showing that even a simple uniform method for balancing the number of instructions and responses can lead to significant improvements. Building on this, we further propose Balanced Preference Optimization (BPO), designed to dynamically augment the knowledge depth of each sample. BPO is motivated by the observation that the usefulness of knowledge varies across samples, necessitating tailored learning of knowledge depth. To achieve this, we introduce gradient-based clustering, estimating the knowledge informativeness and usefulness of each augmented sample based on the model’s optimization direction. Our experimental results across various benchmarks demonstrate that BPO outperforms other baseline methods in alignment tuning while maintaining training efficiency. Furthermore, we conduct a detailed analysis of each component of BPO, providing guidelines for future research in preference data optimization.
@@
Sizhe Wang1  Yongqi Tong2  Hengyuan Zhang3  Dawei Li4  Xin Zhang2  Tianlong Chen5 1University of Southern California, 2Ant Group 3Tsinghua University, 4Arizona State Univeristy, 5University of North Carolina at Chapel Hill sizhewan@usc.edu, {tongyongqi.yq, evan.zx}@antgroup.com, zhang-hy22@mails.tsinghua.edu.cn, daweili5@asu.edu, tianlong@cs.unc.edu
@@ 1 Introduction
Figure 1: Overview of knowledge breadth and depth, and how we link them with the number of instruction and response in alignment tuning datasets.
@@
To address this problem, we begin by proposing the concepts of knowledge breadth and knowledge depth, which measure the comprehensiveness and depth of an LLM or knowledge source respectively. Building on these concepts, we conduct a systematic analysis of the resources allocated for enhancing the breadth and depth of knowledge during alignment tuning accordingly, uncovering a potential imbalance stemming from the structure of alignment datasets. To validate this imbalance, we propose a simple balance method, that includes knowledge breadth compression (KBC) and knowledge depth augmentation (KDA), achieved by uniformly reducing the number of prompts and increasing the number of response pairs in the preference learning dataset. Figure 1 demonstrates the overview process. The substantial improvement brought by both KBC and KDA confirms the effectiveness of our balance theory in alignment tuning datasets.
@@
  • Through extensive experiments, we show that BPO outperforms previous preference data optimization methods. We also conduct further exploration for BPO to provide in-depth hints.
@@ 2.1 Knowledge Breadth and Depth
While LLMs demonstrate striking knowledge breadth in extensive domains and areas, their knowledge depth for providing really in-depth and expert-level output is still not promising Zhang et al. (2024); Bai et al. (2024). In this work, we propose to address this problem by first analyzing the relationship between knowledge breadth and depth, with prompts and responses in the alignment tuning dataset. Intuitively, a large and diverse set of instructions allows the LLMs to cover a wide range of topics and areas, thus expanding its knowledge breadth. Similarly, more responses of various quality provide LLMs with great opportunities to fully understand the question and explore the required knowledge, thus leading to LLMs being able to provide in-depth insights and responses.
However, if we take a closer look at alignment tuning datasets, there is a significant imbalance between the number of instructions and responses: a typical alignment tuning dataset is in the format of $\mathcal{D}=\{(x_{1},y_{1}^{1},y_{1}^{2}),...,(x_{n},y_{n}^{1},y_{n}^{2})\}$, in which each sample ($x,y^{1},y^{2}$) encompasses a prompt $x$, the winning response $y^{1}$ and the losing response $y^{2}$. In this dataset, the instruction number $n$ is usually tens of thousands and the response number is simply 2, where we have $n>>2$. Here we argue that this imbalance between instruction and response numbers actually implies the unbalanced resource allocation for knowledge breadth and depth learning in alignment tuning time, limiting LLMs’ exploration for more in-depth knowledge.
@@ 2.3 Preliminary Experiment & Analysis
In Figure 2, we demonstrate that using embedding-based clustering to select only 1% to 10% of the representative instructions for alignment can achieve performance comparable to, and even surpassing, alignment on the full dataset while significantly reducing computational resource requirements. This finding is consistent with that of Zhou et al. (2024), indicating that data quantity is not the key factor. By focusing on samples at the centroids of clusters, we ensure the quality of the selected prompts while reducing the probability of redundant prompts being selected, thus achieving better performance with fewer prompts. This indicates that a small subset of carefully selected instructions could already satisfy the LLMs expanding in knowledge breadth.
@@ 3.2 Dynamic Knowledge Depth Augmentation
To augment the knowledge depth $\mathcal{K}$ for a given preference learning dataset, we first tune the base model with LoRA Hu et al. (2021) using a subset of seed data and obtian the supervised fine-tuned (SFT) policy $\pi_{\theta}$, more details are in Appendix B. Then, we produce multiple responses for each prompt $x_{i}\in\mathcal{X}_{\text{rep}}$ using the SFT policy $\pi_{\theta}$ and follow previous work to adopt the LLM-as-a-judge Zheng et al. (2023a); Gao et al. (2023); Li et al. (2024) and reject sampling Khaki et al. (2024b) approach to construct high-quality response pairs. Furthermore, since the SFT policy has already been optimized for common scenarios, its responses are typically safe and exhibit minimal differences, making it challenging to generate response pairs with significant score differences. To address this for the safety dataset, we employed a novel method by utilizing jailbreaking prompts to elicit more diverse responses. We analyze this method and its effect in Section 5.3.
@@
More results and analysis about the hyperparameter $\eta\%$ are shown in Table 5. Intuitively, informative augmented samples with gradient vectors close to the centers will get a greater opportunity to be selected, thus being allocated with more depth learning resources. Figure 4 shows the dynamic allocation of response pairs for different prompts, where the number of pairs varies based on the gradient-based clustering and selection. This illustrates how BPO dynamically adjusts the number of response pairs for different samples to optimize the knowledge depth allocation. We then perform DPO fine-tuning using $\mathcal{D}_{dyn}$, which leads to better alignment and performance improvements as demonstrated in our experiments.
@@
Table 1: Performance comparison of different pair selection strategies on the SafeRLHF and HH-RLHF datasets. [1] stands for utilizing Simple Balance on 10% randomly selected samples. As we can see, our method BPO can achieve comparable or even better performance with no more than 10% overall data.
@@ 4.1 Experimental Setup
We choose HH-RLHF Bai et al. (2022) and safeRLHF Dai et al. (2023) as our training set. More details of this two datasets are presented in Appendix G.
@@
We conduct experiment using Llama-2-13B (Touvron et al., 2023) and Llama-3-8B (Dubey et al., 2024). We also utilize Llama3-8b as the feature extractor to obtain representations used for clustering, which serves as a common choice in many clustering work (Petukhova et al., 2024). We have chosen the MT-Bench (Zheng et al., 2023a) and AlpacaEval (Li et al., 2023) to evaluate the effectiveness of our proposed methods and baseline models. We set the scaling ratio for knowledge breadth and depth balance to be 0.1. More evaluation details are attached in Appendix G and the training hyperparameters are shown in Appendix H.
@@ 4.2 Experimental Results
In this section, we compare our overall BPO methods with mentioned baselines and sampling methods. As shown in Table 1, BPO shows comparable or even greatly superior performance with only no more than 10% data. KBC underperforms with BPO as it only relies on using fewer representative instructions, which further demonstrate the necessity for our hierichical sampling approach to balance between both breadth and depth. Compared to simple balance, BPO achieves better performance, which we attribute to the gradient-based dynamic knowledge depth augmentation selecting the most useful and informative samples for model updates. Additionally, potentially redundant samples are less likely to be selected during sampling. This reduces the possible negative impact of knowledge on the model and increases training efficiency. Furthermore, we compared our method with the approach of randomly selecting 10% of the samples used in simple balance. The results show that our gradient clustering performs better, indicating that BPO successfully allocates appropriate knowledge depth resources, thereby enhancing model performance. Furthermore, we compared our method with Curry-DPO and RS-DPO. Notably, RS-DPO’s performance ranks second only to ours, even surpassing BPO on MT-Bench when trained on the HH-RLHF dataset. However, it requires significantly more data than our approach. Overall, our method is both effective and efficient, demonstrating robustness across two datasets and two model sizes while successfully achieving alignment.
@@ 5.1 Clustering Settings
We investigate how the number of clusters affects both embedding vector clustering during the knowledge breadth compression step and gradient clustering in the dynamic knowledge depth augmentation step. Figure 5 shows that for embedding vector clustering and gradient clustering, the results are relatively stable when the number of clusters is 100 and 50, respectively. A cluster number that is too small may fail to adequately represent the comprehensive level of knowledge breadth, making it difficult to achieve stable results and leading to significant fluctuations. Conversely, an excessively large number of clusters might over-segment the data, which does not necessarily guarantee a steady improvement in performance.
@@ 5.2 Measurement of Required Knowledge Depth
Table 2 illustrates that relying solely on response length is inadequate, as it overlooks semantic differences. Responses of similar length can have different meanings, leading to an incomplete exploration of the data. Although semantic similarity performs better than response length, it still underperforms our gradient-based methods. While semantic similarity captures the correlationships within the responses of each prompt, gradient feature represents knowledge breadth and depth on a global scale, resulting in more effective optimization.
@@ 5.3 Comparisons about Augmentation Methods in KDA
As described in Table 3, common response generation methods adopted in many online DPO variants on the safeRLHF dataset surprisingly did not yield good results. Upon analyzing the bad cases, we found that the base model had already been optimized for common scenarios, making its responses quite safe and difficult to obtain unsafe responses.
Therefore, we employed a different method in the safety alignment area: for half of the generations, we used jailbreaking prompts proposed in Shen et al. (2024) to induce the model to produce potentially risky responses. These were then paired with normally generated safe responses for further processing. The pair construction process is the same as described in Section 3.2. Our findings also support the assumption that uniformly increasing the number of more distinctly different pairs between chosen and rejected responses can notably enhance the model’s alignment effects. Moreover, in the context of online RLHF research focused on safety alignment, we want to emphasize to future work the importance of assessing the quality differences in generations. It’s crucial to determine whether the model is already sufficiently safe, as this could hinder the generation of adequately poor negative responses.
@@ 7 Conclusions
Despite the promising results of our hierarchical clustering-based framework, several limitations must be acknowledged. First, learning directly from the model’s generated responses leads to limited or marginal performance improvement, indicating inherent constraints in self-optimization. To achieve high-quality data curation, we rely on the GPT-4 model to score responses to construct response pairs, which introduces a certain level of dependency on the performance and potential biases of GPT-4.
For future work, it is crucial to develop statistical or empirical methods to evaluate whether a user query be beneficial to the entire alignment optimization. Additionally, refining the generation and construction of preference pairs is essential, as it fundamentally shapes the alignment learning space. Further exploration into these areas could yield significant advancements in the alignment capabilities of LLMs.
@@ Appendix H Training Hyperparameters
Appendix I LLM-as-a-Judge’s Prompt
@@
Appendix J Acknowledgment of AI Assistance in Writing and Revision

### v2

@@ Abstract
Reinforcement Learning with Human Feedback (RLHF) is the key to the success of large language models (LLMs) in recent years. In this work, we first introduce the concepts of knowledge breadth and knowledge depth, which measure the comprehensiveness and depth of an LLM or knowledge source, respectively. We reveal that the imbalance in the number of instructions and responses can lead to a potential disparity in breadth and depth learning within alignment tuning datasets by showing that even a simple uniform method for balancing the number of instructions and responses can lead to significant improvements. Building on this, we further propose Balanced Preference Optimization (BPO), designed to dynamically augment the knowledge depth of each sample. BPO is motivated by the observation that the usefulness of knowledge varies across samples, necessitating tailored learning of knowledge depth. To achieve this, we introduce gradient-based clustering, estimating the knowledge informativeness and usefulness of each augmented sample based on the model’s optimization direction. Our experimental results on various benchmarks demonstrate that BPO outperforms other baseline methods in alignment tuning while maintaining training efficiency. Furthermore, we conduct a detailed analysis of each component of BPO, providing guidelines for future research in preference data optimization. Our code will be published soon on https://github.com/Sizhe04/Balanced-Preference-Optimization.
@@
Sizhe Wang1  Yongqi Tong2  Hengyuan Zhang3  Dawei Li4  Xin Zhang2  Tianlong Chen5 1University of Southern California, 2Ant Group, 3Tsinghua University 4Arizona State Univeristy, 5University of North Carolina at Chapel Hill sizhewan@usc.edu, {tongyongqi.yq, evan.zx}@antgroup.com, zhang-hy22@mails.tsinghua.edu.cn, daweili5@asu.edu, tianlong@cs.unc.edu
@@ 1 Introduction
Figure 1: Overview of knowledge breadth and depth, and how we link them with the number of instructions and responses in alignment tuning datasets.
@@
To address this problem, we begin by proposing the concepts of knowledge breadth and knowledge depth, which measure the comprehensiveness and depth of an LLM or knowledge source respectively. Building on these concepts, we conduct a systematic analysis of the resources allocated for enhancing the breadth and depth of knowledge during alignment tuning accordingly, uncovering a potential imbalance stemming from the structure of alignment datasets. To validate this imbalance, we propose a simple balance method, that includes knowledge breadth compression (KBC) and knowledge depth augmentation (KDA), achieved by uniformly reducing the number of prompts and increasing the number of response pairs in the preference learning dataset. Figure 1 demonstrates the overview process. The substantial improvement brought by both KBC and KDA confirms the effectiveness of our balance approach in alignment tuning datasets.
@@
  • Through extensive experiments, we show that BPO outperforms previous preference data optimization methods. We also conduct further exploration on BPO to provide in-depth hints.
@@ 2.1 Knowledge Breadth and Depth
While LLMs demonstrate striking knowledge breadth in extensive domains and areas, their knowledge depth for providing truly in-depth and expert-level output is still not promising Zhang et al. (2024); Bai et al. (2024). In this work, we propose to address this problem by first analyzing the relationship between knowledge breadth and depth, with prompts and responses in the alignment tuning dataset. Intuitively, a large and diverse set of instructions allows the LLMs to cover a wide range of topics and areas, thus expanding its knowledge breadth. Similarly, more responses of various quality provide LLMs with great opportunities to fully understand the question and explore the required knowledge, thus leading to LLMs being able to provide in-depth insights and responses.
However, if we take a closer look at alignment tuning datasets, there is a significant imbalance between the number of instructions and responses: a typical alignment tuning dataset is in the format of $\mathcal{D}=\{(x_{1},y_{1}^{1},y_{1}^{2}),...,(x_{n},y_{n}^{1},y_{n}^{2})\}$, in which each sample ($x,y^{1},y^{2}$) encompasses a prompt $x$, the winning response $y^{1}$ and the losing response $y^{2}$. In this dataset, the instruction number $n$ is usually in the tens of thousands, while the response number is simply 2, where we have $n>>2$. Here we argue that this imbalance between instruction and response numbers actually implies the unbalanced resource allocation for knowledge breadth and depth learning during alignment tuning time, limiting LLMs’ exploration for more in-depth knowledge.
@@ 2.3 Preliminary Experiment & Analysis
In Figure 2, we demonstrate that using embedding-based clustering to select only 1% to 10% of the representative instructions for alignment can achieve performance comparable to, and even surpassing, alignment on the full dataset while significantly reducing computational resource requirements. This finding is consistent with that of Zhou et al. (2024), indicating that data quantity is not the key factor. By focusing on samples at the centroids of clusters, we ensure the quality of the selected prompts while reducing the probability of redundant prompts being selected, thus achieving better performance with fewer prompts. This indicates that a small subset of carefully selected instructions could already satisfy the need of LLMs to expand the breadth of knowledge.
@@ 3.2 Dynamic Knowledge Depth Augmentation
To augment the knowledge depth $\mathcal{K}$ for a given preference learning dataset, we first tune the base model with LoRA Hu et al. (2021) using a subset of seed data and obtian the supervised fine-tuned (SFT) policy $\pi_{\theta}$, more details are in Appendix B. Then, we produce multiple responses for each prompt $x_{i}\in\mathcal{X}_{\text{rep}}$ using the SFT policy $\pi_{\theta}$ and follow previous work to adopt the LLM-as-a-judge Zheng et al. (2023a); Gao et al. (2023); Li et al. (2024a) and reject sampling Khaki et al. (2024b) approach to construct high-quality response pairs. Furthermore, since the SFT policy has already been optimized for common scenarios, its responses are typically safe and exhibit minimal diversity, making it challenging to generate response pairs with significant score differences. To address this for the safety dataset, we employed a novel method by utilizing jailbreaking prompts to elicit more diverse responses. We analyze this method and its effects in Section 5.3.
@@
More results and analysis about the hyperparameter $\eta\%$ are shown in Table 5. Intuitively, informative augmented samples with gradient vectors close to the centers will get a greater opportunity to be selected, thus being allocated with more depth learning resources. Figure 4 shows the dynamic allocation of response pairs for different prompts, where the number of pairs varies based on the gradient-based clustering and selection. This illustrates how BPO dynamically adjusts the number of response pairs for different samples to optimize the knowledge depth allocation. We then perform DPO fine-tuning using $\mathcal{D}_{dyn}$, leading to better alignment and performance improvements as demonstrated in our experiments.
@@
Table 1: Performance comparison of different pair selection strategies on the SafeRLHF and HH-RLHF datasets. [1] stands for utilizing Simple Balance on 10% randomly selected samples. As shown, our method BPO can achieve comparable or even better performance with no more than 10% overall data.
@@ 4.1 Experimental Setup
We choose HH-RLHF Bai et al. (2022) and safeRLHF Dai et al. (2023) as our training sets. HH-RLHF consists of two subsets, with approximately 170K chosen-rejected pairs related to human and AI assistant dialogues. For our experiment, we randomly select 30,000 samples from the helpfulness subset. SafeRLHF contains 83.4K preference entries, where each entry includes two responses with labels for harmlessness and helpfulness. In our experiment, we select samples where both labels are the same, resulting in approximately 55,000 samples. Additionally, for each dataset, we randomly select 10% of the dataset $\mathcal{D}_{org}$ as the seed data $\mathcal{D}_{seed}$, using annotated chosen responses as golden reference to perform preliminary SFT.
@@
We conduct experiments using Llama-2-13B (Touvron et al., 2023) and Llama-3-8B (Dubey et al., 2024). We also utilize Llama3-8b as the feature extractor to obtain representations used for clustering, which serves as a common choice in many clustering work (Petukhova et al., 2024). We adopt the MT-Bench (Zheng et al., 2023a) and AlpacaEval (Li et al., 2023) to evaluate the effectiveness of our proposed methods and baseline models. We set the scaling ratio for knowledge breadth and depth balance to 0.1. More evaluation details are attached in Appendix G and the training hyperparameters are shown in Appendix H. Additionally, we conducted further experiments on other datasets to validate the effectiveness and robustness of our method, which are detailed in Appendix I.
@@ 4.2 Experimental Results
In this section, we compare our overall BPO methods with the mentioned baselines and sampling methods. As shown in Table 1, BPO BPO achieves comparable or even superior performance while using no more than 10% of the data. KBC underperforms compared to BPO as it only relies on using fewer representative instructions, which further demonstrate the necessity for our hierichical sampling approach to balance between both breadth and depth. Compared to simple balance, BPO achieves better performance, which we attribute to the gradient-based dynamic knowledge depth augmentation selecting the most useful and informative samples for model updates. Additionally, potentially redundant samples are less likely to be selected during sampling. This reduces the possible negative impact of knowledge on the model and increases training efficiency. Furthermore, we compared our method with the approach of randomly selecting 10% of the samples used in simple balance. The results show that our gradient clustering performs better, indicating that BPO successfully allocates appropriate knowledge depth resources, thereby enhancing model performance. Furthermore, we compared our method with Curry-DPO and RS-DPO. Notably, RS-DPO’s performance ranks second only to ours, even surpassing BPO on MT-Bench when trained on the HH-RLHF dataset. However, it requires significantly more data than our approach. Overall, our method is both effective and efficient, demonstrating robustness across two datasets and two model sizes while successfully achieving alignment.
@@ 5.1 Clustering Settings
We investigate how the number of clusters affects both embedding vector clustering during the knowledge breadth compression step and gradient clustering in the dynamic knowledge depth augmentation step. Figure 5 shows that when using 100 clusters for embedding vector clustering and 50 clusters for gradient clustering, the results are relatively stable. A cluster number that is too small may fail to adequately represent the comprehensive level of knowledge breadth, making it difficult to achieve stable results and leading to significant fluctuations. Conversely, an excessively large number of clusters might over-segment the data, which does not necessarily guarantee a steady improvement in performance.
@@ 5.2 Measurement of Required Knowledge Depth
Table 2 illustrates that relying solely on response length is inadequate, as it overlooks semantic differences. Responses of similar length can have different meanings, leading to an incomplete exploration of the data. Although semantic similarity performs better than response length, it still underperforms our gradient-based methods. While semantic similarity captures the correlations within the responses of each prompt, gradient feature represents knowledge breadth and depth on a global scale, resulting in more effective optimization.
@@ 5.3 Comparisons about Augmentation Methods in KDA
As described in Table 3, common response generation methods adopted in many online DPO variants on the safeRLHF dataset surprisingly did not yield promising results. Upon analyzing the bad cases, we found that the base model had already been optimized for common scenarios, making its responses quite safe and difficult to obtain unsafe ones.
Therefore, we employed a different method in the safety alignment area: for half of the generations, we used jailbreaking prompts proposed in Shen et al. (2024) to induce the model to produce potentially risky responses. These were then paired with normally generated safe responses for further processing. The pair construction process is the same as described in Section 3.2. Our findings also support the assumption that uniformly increasing the number of more distinctly different pairs between chosen and rejected responses can notably enhance the model alignment effects. Moreover, in the context of online RLHF research focused on safety alignment, we want to emphasize to future research the importance of assessing quality differences in generations. It is crucial to determine whether the model is already sufficiently safe, as this could hinder the generation of adequately poor negative responses.
@@ 7 Conclusions
Despite the promising results of our hierarchical clustering-based framework, several limitations must be acknowledged. First, learning directly from the model’s generated responses leads to limited or marginal performance improvement, indicating inherent constraints in self-optimization. To achieve high-quality data curation, we rely on the GPT-4 model to score responses to construct response pairs, which introduces a certain level of dependence on the performance and potential biases of GPT-4.
For future work, it is crucial to develop statistical or empirical methods to evaluate whether a user query would be beneficial to the entire alignment pipeline. Additionally, refining the generation and construction of preference pairs is essential, as it fundamentally shapes the alignment learning space. Further exploration into these areas could yield significant advancements in the alignment capabilities of LLMs.
@@ Appendix H Training Hyperparameters
Appendix I Additional Experiments
To further demonstrate that BPO is both effective and robust, we conducteded additional experiments on the UltraFeedback dataset Cui et al. (2023), anselectedcted AlpacaEval 2.0 Dubois et al. (2024) and Arena-Hard Li et al. (2024b) as our evaluation benchmarks. All experiments were carried out using the Llama-3-8B-Instruct model.
Trainset
UltraFeedback contains AI-generated feedback annotations for a wide variety of prompts, facilitating preference modeling. For our experiments, we randomly select 30,000 samples from UltraFeedback, of which 10% are used as seed data $\mathcal{D}_{seed}$, following the setup described in Section 4.1. We used the same training hyperparameters as presented in the Appendix H.
Evaluation Set
AlpacaEval 2.0 comprises 805 questions drawn from five different datasets, while Arena-Hard includes 500 well-defined technical problem-solving queries. We report all scores according to the respective evaluation protocols of the benchmarks. For AlpacaEval 2.0, we report both the raw win rate and the length-controlled win rate (LC), where the LC metric is designed to be robust against model verbosity. For Arena-Hard, we report the win rate compared to the baseline model.
Method AlpacaEval 2.0 Arena Hard
LC Win Rate (%) Win Rate (%) Win Rate (%)
Base-model 22.92 22.57 20.20
Vanilla DPO 28.71 27.24 33.21
Curry-DPO 35.46 36.65 34.71
RS-DPO 41.23 40.44 37.98
BPO (ours) 44.75 43.68 39.83
Table 6: Performance comparison of different pair selection strategies on the UltraFeedback dataset.
As shown in Table 6, our BPO method demonstrates substantial improvements over the base model and other baselines on both the AlpacaEval 2.0 and Arena Hard datasets. These additional experiments further validate the robustness and effectiveness of BPO across diverse datasets and models.
Appendix J LLM-as-a-Judge’s Prompt
@@ Appendix I Additional Experiments
Appendix K Acknowledgment of AI Assistance in Writing and Revision